# o Mata-Cervera0000-0001-8464-5102 on N. Vetlugin0000-0002-2480-0462 are Soci0000-0002-0149-9128 uel A. Porras0000-0001-8058-9377 ie Shen0000-0002-6700-9902

### v1

@@ Abstract
Identifying physical mechanisms that protect the information carried by various forms of structured light is one of the cornerstones of today’s classical and quantum communications. Here we show that the purity of orbital angular momentum (OAM) modes can be protected against degradation in random media by leveraging two fundamental features of their own Schrödinger Hamiltonian dynamics, namely, Zeno effect —frequent observations slow down the evolution—, and Gouy phase —the back-action of the observation. Repeated, OAM-dependent Gouy phase kicks imparted along the disturbing path by simple imaging systems trigger the optical Zeno effect that protects the input OAM mode against mode cross-talk that would broaden the OAM spectrum. Given the universality of the mechanism, the Gouy phase-assisted Zeno effect would protect propagation modes other than those of OAM, and the diverse forms of structured light built with them.
@@
The simplest form of wave structuring is spatial localization. Then, its propagation phase gets reduced compared to that of a plane wave, a phenomenon known as the Gouy phase anomaly [gouy1890propriete, gouy1890propagation]. It was first described in optics, but affects all waves whatever their nature —quantum, acoustic, or electromagnetic [holme2002gouy, ducharme2015gouy, hiekkamaki2022observation, hamazaki2006direct]—, although its ultimate origin remains a matter of debate [subbarao1995topological, boyd1980intuitive, feng2001physical]. Gouy phase causes diverse phenomena such as the rotation of the energy flow vector [padgett1995poynting] and the polarization state [chen2025gouy], subluminal group velocity [bareza2016subluminal, giovannini2015spatially], longitudinal synchronization and pattern revivals [da2020pattern, zhong2021gouy], to name a few. Gouy phase gradual reduction through a focus differs in different wave models and for particular waves [forbes2021structured, martelli2010gouy, volyar1999topological, cruz2020laguerre, feng1998gouy].
For paraxial light beams obeying the free Schrödinger equation, Gouy phase follows the renowned inverse tangent law of variation for the fundamental mode of propagation, the Gaussian beam, and multiples of it for higher-order modes such as Hermite-Gauss and Laguerre-Gauss (LG) beams. This remarkable mode-dependence of Gouy phase has been exploited to sort different LG modes of a structured beam [gu2018gouy].
Another fundamental feature of the Schrödinger dynamics, more acknowledged in quantum physics, is the short-time parabolic decay of the survival probability of an initial quantum state, a feature that makes possible the Zeno effect (ZE). The evolution of the state can be slowed down or even halted through frequent monitoring of whether the state remains the initial one. Initially proposed by Alan Turin [Turing] as a quantum paradox, then proved by Misra et al. [zeno_misra_1], and later generalized to quantum Zeno dynamics by Facchi et al. [facchi2000quantum], the ZE has inspired a variety of quantum Zeno evolutions [zeno_computing_1, QZD_experimental, barontini2025quantum, sazonov2023quantum]. The original projective measurements in the ZE are often replaced by other models of measurement, such as interaction with other systems (the measuring apparatus) that break the coherence between the system states. Beyond measurements, it is now recognized that other frequent interruptions of the Hamiltonian evolution that alter the relative phases between states, such as unitary kicks, also produce ZE. [QZD_experimental, zeno_integrated_photonics, blanchard1993strongly, three_examples_QZD, QZD_review].
@@
The ZE has a classical analog that has been efficiently implemented to modify the evolution of classical light. For example, segmented arrays of coupled waveguides freeze classical tunneling decay from the initial waveguide [longhi_waveguides], or disconnect tunneling-coupled waveguides [zeno_integrated_photonics]. Frequent repeaters restoring an image in a communication channel protect it from noise degradation [gu2012repeater]. In free-space propagation of paraxial light beams, frequent diffracting slits suppress diffraction [diffraction_spreading] and the increasing coherence on propagation [partially_coherent].
Gouy phase is revealed here as a potential tool for Zeno-protection of structured light from degradation in random media, such as atmospheric turbulence [malik2012influence]. The degradation effect of turbulence on a single input mode of orbital angular momentum (OAM), or LG beam, manifests itself as a mode cross-talk that transfers power to other OAM modes, broadening the OAM spectrum [structlightturb, klug2021orbital, bachmann2024universal], and as a decay of the fraction of power in the initial OAM mode —the survival probability of the initial state in quantum physics. In this Letter we show that these effects are strongly suppressed when the turbulent propagation is frequently interrupted by Gouy phase kicks. These kicks introduce an OAM-dependent Gouy phase that would be acquired on free space propagation, which in practice are imparted by adequate optics, such as the $4f$ imaging system in the conceptual Fig. 1. We demonstrate that these repeated Gouy phase kicks produce a ZE that strongly inhibits mode cross-talk and protects the purity of the OAM spectrum.
@@
We first calculate the refractive index $\Delta n(x,y,z)$ within the whole volume of propagation, and then Eq. (1) is solved using a standard split-step numerical algorithm [turb_mask_1, turb_mask_2]. Each refractive index distribution $\Delta n(x,y,z)$ models a random realization of a turbulent medium described by Kolmogorov’s theory [kolmogorov1991local, imaging_turbulence, turb_mask_2] (see details in Annex (I.2)). To account for the stochastic nature of turbulence, we model light propagation in multiple “trials”, each time performing an independent random realization of the refractive index volume. Then, we perform statistics of the results by taking mean values and standard deviations (this method is used throughout the paper) [klug2021orbital, structlightturb].
As an example, we consider propagation through the turbulent medium of a LG mode $\psi_{\ell,p}(\bm{r})$ with $\left(\ell,p\right)=\left(1,0\right)$, where $\ell$ and $p$ stand for azimuthal and radial numbers [see the explicit expression in Annex (I.1)]. Optical beams that propagate in free space as eigenmodes (e.g., LG modes) no longer preserve this property in complex media. Indeed, in our example, the $\left(\ell,p\right)=\left(1,0\right)$ mode, $\psi(\bm{r}_{\perp},0)=\psi_{1,0}(\bm{r}_{\perp},0)$, prepared at $z=0$ evolves into a superposition of LG modes,
@@
where the propagation-dependent complex coefficients are given by the overlap integral $c_{l,p}(z)=\iint\psi_{l,p}^{\star}(\bm{r}_{\perp},z)\psi(\bm{r}_{\perp},z)\textrm{d}\bm{r}_{\perp}$. Equation (3) describes mode cross-talk [bachmann2024universal, peters2025structured] that decreases the survival power (SP),
@@
the final SP after all measurements is significantly enhanced, as confirmed by modeling in Fig. 2(c). The lower blue curve shows the initial mode’s SP decay during uninterrupted propagation, with a single measurement performed at the end ($N=1$). Following Zeno’s dichotomy paradox [aristotle], the distance between the successive measurements is progressively halved: each subsequent curve represents a modeling of propagation through the same medium, but interrupted by projective measurements every half ($N=2$), quarter ($N=4$), eighth ($N=8$), and so on, of the total propagation distance $z_{\rm max}$. As $N$ increases, mode cross-talk is progressively reduced, leading to higher SP at output, as explicitly shown in Fig. 2(d). This is the hallmark of the Zeno effect. In addition, the OAM-carrying light is getting less susceptible to the randomness of turbulent medium as the decreasing standard deviation in Fig. 2(d) confirms.
Unfortunately, projective measurements of single LG modes lack practical physical implementation [gagliardi1976optical, gibson2004free, yan2014high]. Moreover, this approach is incompatible with communication protocols since it requires projection onto a predetermined mode and complicates the implementation of OAM encoding and information transfer. Nevertheless, the Zeno effect can be achieved by other means, including decoherence, strong coupling and unitary kicks [zeno_integrated_photonics, QZD_experimental, blanchard1993strongly, three_examples_QZD, QZD_review]. Below we show that the ZE driven by unitary kicks can indeed be induced in OAM modes propagating through turbulence by exploiting the mode-dependent nature of the Gouy phase shifts in easily implementable systems that are also compatible with communication protocols.
@@
with $\Phi=\pi$ producing the strongest “decoupling” effect [dhar2006preserving, facchi2004unification, facchi2019kick]. Frequent interruption of propagation through turbulence by unitary kicks reduces the probability of cross-talk of the initial mode with other modes. As shown in Fig. 3(a), frequent $\hat{U}_{k}$ results in an even more pronounced slowdown of the SP decay than with projections. Nonetheless, as happened with projections (5), we still wonder what optical operation yields $\hat{U}_{k}$ (7), i.e., attributes different phases to the mode to be protected and to the cross-talk terms.
Among all mode-dependent properties of the LG beams in (A1), Gouy phase, $\mathcal{G}_{lp}(z)=-(|\ell|+2p+1)\tan^{-1}(z/z_{R})$, where $z_{R}$ is the Rayleigh distance, is particularly relevant as each mode acquires a different phase on free space propagation. In practice, we do not need free space propagation, but a portion or the complete Gouy phase shift from $\left(-\infty,\infty\right)$ can be imparted by simple optics.
For example, when a LG beam passes through a $2f$-focusing system ($\hat{U}_{k}^{2f}$) with $f=z_{R}$, it is transformed into $\hat{U}_{k}^{2f}\psi_{\ell p}(\bm{r}_{\perp})=\textrm{exp}\left(i\left[1+|\ell|+2p\right]\pi/2\right)\psi_{\ell p}(\bm{r}_{\perp})$, provided that $\psi_{\ell p}(\bm{r}_{\perp})$ is collimated (negligible wavefront curvature). The condition $f=z_{R}$ just makes the input and output mode components have the same size, thus, it only experiences a phase shift. An arbitrary input beam (3) in the $2f$-system will be transformed into
@@
There is a significant difference between projections $\hat{\Pi}$ and kicks $\hat{U}_{k}$: each individual $\hat{\Pi}$ eliminates all cross-talk terms, effectively “cleaning” the beam and restarting the evolution. The action of individual $\hat{U}_{k}$ neither alters the power in each mode, $P_{\ell p}=|c_{\ell p}|^{2}$, nor the OAM spectrum and its expected value $\langle L_{z}\rangle=\sum_{\ell p}\ell\cdot|c_{\ell p}|^{2}$. It is the repeated action of multiple $\hat{U}_{k}$ that affects the whole evolution, reducing the mode cross-talk spread of the OAM spectrum. Neither $\hat{U}_{k}^{2f}$ nor $\hat{U}_{k}^{4f}$ are $\hat{U}_{k}$ as defined in (7), but they also produce ZE.
@@
Finally, we discuss the existence of ZE in stronger turbulent conditions. A key parameter that determines the turbulence strength is the coherence length $r_{0}$ (B2) which measures the scale at which the medium’s refractive index becomes uncorrelated [structlightturb, bachmann2024universal]. The distortion experienced by the beam will be ruled by the ratio $\langle r\rangle/r_{0}$ between the beam size $\langle r\rangle$ and the coherence length, since larger beams experience more randomness than smaller beams in the same medium [klug2021orbital, bachmann2024universal]. Here we wonder if ZE will still occur if the disturbance experienced by the beam increases, that is, for higher values of $w_{0}/r_{0}$.
@@
Beams with larger beam waist $w_{0}$ present faster SP decays as shown in Fig. 4(a). When we introduce an increasing number of $\pi-\hat{U}_{k}$, the decay of the SP is slowed down, and this seems to happen for any input beam size $w_{0}$ as shown in (b). It can be observed in previous figures 2(b) and 3(d) that there is a certain range in which increasing $N$ does not increase significantly the SP at the final plane, but once $N$ exceeds a certain value the power increases much faster. The same behavior is found for $\pi-\hat{U}_{k}$’s in Fig. 4(c), where an abrupt increase of the SP happens once the distance between kicks decreases and becomes comparable with the coherence length $r_{0}$. This phenomenon is common to all simulated values of $w_{0}$ but it becomes more obvious when the decay rate is faster, for larger values of $w_{0}$. As shown in 4(d), the larger $w_{0}$ the lower the SP at the final plane without introducing any $\hat{U}_{k}$ due to a faster decay rate. At first, the SP increases only very slightly when we increase the frequency of $\hat{U}_{k}$ - weak ZE. When the distance between $\hat{U}_{k}$ $(\Delta z_{k})$ becomes of the order of the coherence length $r_{0}$ or smaller the SP increases significantly for larger $N$: strong ZE. This distinguishes two regimes: low interruption frequency $(\Delta z_{k}\gg r_{0}$, green shaded area in Fig.  4(d)) in which turbulence effects dominate, and the Zeno regime $(\Delta z_{k}\lesssim r_{0}$, red shaded area in Fig.  4(d)), in which increasing $N$ reduces mode cross-talk significantly. The transition between these two regimes happens when $\Delta z_{k}\sim r_{0}$ for any beam size, suggesting a connection between the coherence length $r_{0}$ and the Zeno distance – the distance between measurements that yields ZE, a spatial analog of the Zeno time in quantum systems [three_examples_QZD, nakazato1996temporal, wilkinson1997experimental, anastopoulos2019decays]. One would expect the Zeno distance to decrease for increasing randomness (higher $C_{n}^{2}$), requiring higher interruption rate to maintain a certain OAM purity, which could make this approach unpractical or inefficient in severe turbulence conditions. On the other hand, if we fix the number of interruptions, their ability to produce Zeno effect will improve as the Zeno distance increases. Given that the results in Fig. (2,3,4) were obtained with very severe turbulence (the survival power decays within $4.5\text{\,}\mathrm{m}$), an additional simulation is carried out scaling the evolution to more realistic distances of the order of hundreds of meters, as shown in the Annex section (I.3). In this simulation, the distance between consecutive $\hat{U}_{k}^{2f}$ increases to the order of tens of centimetres. The feasibility to implement this approach in realistic channels will depend on the turbulence strength: while moderate or weak turbulence requires reasonable distances between interruptions, the required distance for stronger turbulence may become considerably small. This limitation may be overcome by replacing the proposed lens-based Gouy phase kicks by more practical mechanisms.

### v2

@@ Abstract
Identifying physical mechanisms that preserve the integrity of structured waves under propagation through complex environments is one of the central problems in wave physics. Here we show that the purity of orbital angular momentum (OAM) modes can be protected against degradation in random media by leveraging two fundamental features of their own Schrödinger Hamiltonian dynamics, namely, Zeno effect —frequent observations slow down the evolution—, and Gouy phase —the back-action of the observation. Repeated, OAM-dependent Gouy phase kicks imparted along the disturbing path by simple imaging systems trigger the optical Zeno effect that protects the input OAM mode against mode cross-talk that would broaden the OAM spectrum. Given the universality of the mechanism, the Gouy phase-assisted Zeno effect would protect propagation modes other than those of OAM, and the diverse forms of structured light built with them.
@@
I Introduction
The simplest form of wave structuring is spatial localization. Then, its propagation phase gets reduced compared to that of a plane wave, a phenomenon known as the Gouy phase anomaly [25, 26]. It was first described in optics, but affects all waves whatever their nature —quantum, acoustic, or electromagnetic [32, 14, 30, 29]—, although its ultimate origin remains a matter of debate [49, 8, 20]. Gouy phase causes diverse phenomena such as the rotation of the energy flow vector [42] and the polarization state [9], subluminal group velocity [4, 24], longitudinal synchronization and pattern revivals [12, 54], to name a few. Gouy phase gradual reduction through a focus differs in different wave models and for particular waves [21, 39, 51, 11, 19].
For paraxial light beams obeying the free Schrödinger equation, Gouy phase follows the renowned inverse tangent law of variation for the fundamental mode of propagation, the Gaussian beam, and multiples of it for higher-order modes such as Hermite-Gauss and Laguerre-Gauss (LG) beams. This remarkable mode-dependence of Gouy phase has been exploited to sort different LG modes of a structured beam [28].
Another fundamental feature of the Schrödinger dynamics, more acknowledged in quantum physics, is the short-time parabolic decay of the survival probability of an initial quantum state, a feature that makes possible the Zeno effect (ZE). The evolution of the state can be slowed down or even halted through frequent monitoring of whether the state remains the initial one [31, 40, 15]. The original projective measurements in the ZE are often replaced by other models of measurement, such as interaction with other systems (the measuring apparatus) that break the coherence between the system states. Beyond measurements, it is now recognized that other frequent interruptions of the Hamiltonian evolution that alter the relative phases between states, such as unitary kicks, also produce ZE. [48, 37, 7, 5, 17].
@@ I Introduction
The ZE has a classical analog that has been efficiently implemented to modify the evolution of classical light. For example, segmented arrays of coupled waveguides freeze classical tunneling decay from the initial waveguide [6], or disconnect tunneling-coupled waveguides [37]. Frequent repeaters restoring an image in a communication channel protect it from noise degradation [27]. In free-space propagation of paraxial light beams, frequent diffracting slits suppress diffraction [44] and the increasing coherence on propagation [45].
Gouy phase is revealed here as a theoretical tool for Zeno-protection of structured light from degradation in random media, such as atmospheric turbulence [38]. The degradation effect of turbulence on a single input mode of orbital angular momentum (OAM), or LG beam, manifests itself as a mode cross-talk that transfers power to other OAM modes, broadening the OAM spectrum [10, 33, 3], and as a decay of the fraction of power in the initial OAM mode —the survival probability of the initial state in quantum physics. In this paper we show that these effects are strongly suppressed when the turbulent propagation is frequently interrupted by Gouy phase kicks. These kicks introduce an OAM-dependent Gouy phase that would be acquired on free space propagation, which in practice are imparted by adequate optics, such as the $4f$ imaging system in the conceptual Fig. 1. We demonstrate that these repeated Gouy phase kicks produce a ZE that strongly inhibits mode cross-talk and protects the purity of the OAM spectrum.
II Light propagation in turbulent media
@@
We first calculate the refractive index $\Delta n(x,y,z)$ within the whole volume of propagation, and then Eq. (1) is solved using a standard split-step numerical algorithm [46, 36]. Each refractive index distribution $\Delta n(x,y,z)$ models a random realization of a turbulent medium described by Kolmogorov’s theory [35, 47, 36] [see details in Appendix (VI.2)]. To account for the stochastic nature of turbulence, we model light propagation in multiple “trials”, each time performing an independent random realization of the refractive index volume. Then, we perform statistics of the results by taking mean values and standard deviations (this method is used throughout the paper) [33, 10].
As an example, we consider propagation through the turbulent medium of a LG mode $\psi_{\ell,p}(\bm{r})$ with $\left(\ell,p\right)=\left(1,0\right)$, where $\ell$ and $p$ stand for azimuthal and radial numbers [see the explicit expression in Appendix (VI.1)]. Optical beams that propagate in free space as eigenmodes (e.g., LG modes) no longer preserve this property in complex media. Indeed, in our example, the $\left(\ell,p\right)=\left(1,0\right)$ mode, $\psi(\bm{r}_{\perp},0)=\psi_{1,0}(\bm{r}_{\perp},0)$, prepared at $z=0$ evolves into a superposition of LG modes,
@@
where the propagation-dependent complex coefficients are given by the overlap integral $c_{l,p}(z)=\iint\psi_{l,p}^{\star}(\bm{r}_{\perp},z)\psi(\bm{r}_{\perp},z)\textrm{d}\bm{r}_{\perp}$. Equation (3) describes mode cross-talk [3, 43] that decreases the survival power (SP),
@@ III.1 Projective measurements
the final SP after all measurements is significantly enhanced, as confirmed by modeling in Fig. 2(c). The lower blue curve shows the initial mode’s SP decay during uninterrupted propagation, with a single measurement performed at the end ($N=1$). Following Zeno’s dichotomy paradox [2], the distance between the successive measurements is progressively halved: each subsequent curve represents a modeling of propagation through the same medium, but interrupted by projective measurements every half ($N=2$), quarter ($N=4$), eighth ($N=8$), and so on, of the total propagation distance $z_{\rm max}$. As $N$ increases, mode cross-talk is progressively reduced, leading to higher SP at output, as explicitly shown in Fig. 2(d). This is the hallmark of the Zeno effect. In addition, the OAM-carrying light is getting less susceptible to the randomness of turbulent medium as the decreasing standard deviation in Fig. 2(d) confirms.
Unfortunately, projective measurements of single LG modes lack practical physical implementation [22, 23, 53], are inherently lossy and require a projection onto a single predetermined mode. Nevertheless, the Zeno effect can be achieved by other means, including decoherence, strong coupling and unitary kicks [37, 48, 7, 5, 17]. Below we show that the ZE driven by unitary kicks can indeed be induced in OAM modes propagating through turbulence by exploiting the mode-dependent nature of the Gouy phase shifts. Furthermore, we show that these unitary kicks could be in principle implemented using elementary optics such as imaging systems.
III.2 Unitary kicks
III.2.1 Mode-selective unitary kicks
@@
with $\Phi=\pi$ producing the strongest “decoupling” effect [13, 16, 18]. Frequent interruption of propagation through turbulence by unitary kicks reduces the probability of cross-talk of the initial mode with other modes. As shown in Fig. 3(a), frequent $\hat{U}_{k}$ results in an even more pronounced slowdown of the SP decay than with projections. Nonetheless, as happened with projections (5), we still wonder what optical operation yields $\hat{U}_{k}$ (7), i.e., attributes different phases to the mode to be protected and to the cross-talk terms.
III.2.2 Gouy phase-induced unitary kicks
Among all mode-dependent properties of the LG beams in (A1), Gouy phase, $\mathcal{G}_{lp}(z)=-(|\ell|+2p+1)\tan^{-1}(z/z_{R})$, where $z_{R}$ is the Rayleigh distance, is particularly relevant as each mode acquires a different phase on free space propagation. In practice, we do not need free space propagation, but a portion or the complete Gouy phase shift from $\left(-\infty,\infty\right)$ can be imparted by simple optics. For example, when a LG beam passes through a $2f$-focusing system ($\hat{U}_{k}^{2f}$) with $f=z_{R}$, it is transformed into $\hat{U}_{k}^{2f}\psi_{\ell p}(\bm{r}_{\perp})=\textrm{exp}\left(i\left[1+|\ell|+2p\right]\pi/2\right)\psi_{\ell p}(\bm{r}_{\perp})$, provided that $\psi_{\ell p}(\bm{r}_{\perp})$ is collimated (negligible wavefront curvature). The condition $f=z_{R}$ just makes the input and output mode components have the same size, thus, it only experiences a phase shift. An arbitrary input beam (3) in the $2f$-system will be transformed into
@@ III.2.2 Gouy phase-induced unitary kicks
There is a significant difference between projections $\hat{\Pi}$ and kicks $\hat{U}_{k}$: each individual $\hat{\Pi}$ eliminates all cross-talk terms, effectively “cleaning” the beam and restarting the evolution. The action of individual $\hat{U}_{k}$ neither alters the power in each mode, $P_{\ell p}=|c_{\ell p}|^{2}$, nor the OAM spectrum and its expected value $\langle L_{z}\rangle=\sum_{\ell p}\ell\cdot|c_{\ell p}|^{2}$. It is the repeated action of multiple $\hat{U}_{k}$ that affects the whole evolution, reducing the mode cross-talk spread of the OAM spectrum. Notice that neither $\hat{U}_{k}^{2f}$ nor $\hat{U}_{k}^{4f}$ are $\hat{U}_{k}$ as defined in (7), but they also produce (a less efficient) ZE.
@@ IV Effect of turbulence strength and Zeno distance
Finally, we discuss the existence of ZE in stronger turbulent conditions. A key parameter that determines the turbulence strength is the coherence length $r_{0}$ (B2) which measures the scale at which the medium’s refractive index becomes uncorrelated [10, 3]. The distortion experienced by the beam will be determined by the ratio $\langle r\rangle/r_{0}$ between the beam size $\langle r\rangle$ and the coherence length, since larger beams experience more randomness than smaller beams in the same medium [33, 3]. Here we wonder if ZE will still occur if the disturbance experienced by the beam increases, that is, for higher values of $w_{0}/r_{0}$.
@@
Beams with larger beam waist $w_{0}$ present faster SP decays as shown in Fig. 4(a). When we introduce an increasing number of $\pi-\hat{U}_{k}$, the decay of the SP is slowed down, and this phenomenon is observed for any input beam size $w_{0}$ as shown in (b). It can be seen in previous figures 2(b) and 3(d) that there is a certain range in which increasing $N$ does not increase significantly the SP at the final plane, but once $N$ exceeds a certain value the power increases much faster. The same behavior is found for $\pi-\hat{U}_{k}$’s in Fig. 4(c), where an abrupt increase of the SP happens once the distance between kicks decreases and becomes comparable with the coherence length $r_{0}$. This phenomenon is common to all simulated values of $w_{0}$ but it becomes more obvious when the decay rate is faster, for larger values of $w_{0}$. As shown in 4(d), the larger $w_{0}$ the lower the SP at the final plane without introducing any $\hat{U}_{k}$ due to a faster decay rate. At first, the SP increases only very slightly when we increase the frequency of $\hat{U}_{k}$ - weak ZE. When the distance between $\hat{U}_{k}$ $(\Delta z_{k})$ becomes of the order of the coherence length $r_{0}$ or smaller the SP increases significantly for larger $N$: strong ZE. This distinguishes two regimes: low interruption frequency $(\Delta z_{k}\gg r_{0}$, green shaded area in Fig.  4(d)) in which turbulence effects dominate, and the Zeno regime $(\Delta z_{k}\lesssim r_{0}$, red shaded area in Fig.  4(d)), in which increasing $N$ reduces mode cross-talk significantly. The transition between these two regimes happens when $\Delta z_{k}\sim r_{0}$ for any beam size, suggesting a connection between the coherence length $r_{0}$ and the Zeno distance – the distance between measurements that yields ZE, a spatial analog of the Zeno time in quantum systems [5, 41, 52, 1]. One would expect the Zeno distance to decrease for increasing randomness (higher $C_{n}^{2}$), requiring higher interruption rate to maintain a certain OAM purity, which could makes this approach highly unpractical in severe turbulence conditions. On the other hand, if we fix the number of interruptions, their ability to produce Zeno effect will improve as the Zeno distance increases. Given that the results in Fig. (2,3,4) were obtained with very severe turbulence (the survival power decays within $4.5\text{\,}\mathrm{m}$), an additional simulation is carried out scaling the evolution to more realistic distances of the order of hundreds of meters, as shown in the Appendix section (VI.3). In this simulation, the distance between consecutive $\hat{U}_{k}^{2f}$ increases to the order of tens of centimetres. Nonetheless, even in such moderate turbulence scenario the number of lenses needed to produce an observable Gouy phase-assisted Zeno effect is undesirably large, therefore more practical and efficient decoupling mechanisms may be explored in future research.
V Conclusions